OpenStudy (anonymous):

5[1/2+ (3/5 x 5/6) divide 5/8], just checking, is it 7.5??

OpenStudy (anonymous):

uh oh... mine is 6.5 =/

OpenStudy (lalaly):

\[5( \frac{1}{2}+(\frac{15}{30})\div \frac{5}{8})\]

OpenStudy (lalaly):

\[5(\frac{1}{2}+\frac{1}{2} \times \frac{8}{5})\]

OpenStudy (lalaly):

\[5(\frac{1}{2}+\frac{8}{10})\]

OpenStudy (lalaly):

\[5(\frac{5}{10}+\frac{8}{10})\]

OpenStudy (lalaly):

\[5(\frac{13}{10})\]

OpenStudy (lalaly):

\[\frac{13}{2}\]

OpenStudy (lalaly):

\[=6.5\]

OpenStudy (anonymous):

OHH, Thank yoou

OpenStudy (lalaly):

ur welcome :D